OpenStudy (anonymous):

Prove using mathematical induction that 3 divides n(n+1)(n+2). I have the first step for n=1, and then assume 3 divides k(k+1)(k+2), k being an integer. I'm now at the next step of trying to prove 3 divides (k+1)(k+2)(k+3) and am stuck.

OpenStudy (amistre64):

how did you prove the first part :)

OpenStudy (amistre64):

or are we just to assume the k(k+1)(k+2) is divisible by 3

OpenStudy (anonymous):

I have 1(1+1)(1+2)=6 which divides into 3. And then k=n, for k(k+1)(k+2). The examples from my prof go from showing n=1 to assuming the statement for k, then plugging in k+1 for k and solving from there...

OpenStudy (precal):

Assume n=k is valid and sub into n=k+1

OpenStudy (amistre64):

if (k+1)(k+2)(k+3) is a scalar of k(k+1)(k+2); then id say it goes

OpenStudy (anonymous):

how do I prove that it is a scalar?

OpenStudy (amistre64):

dunno, havent done any induced proofs k+1 k+2 ---- k(k^2 +3k +2) (k+1)(k^2 +3k +2) perhaps?

OpenStudy (amistre64):

err, (k+3)(k^2 +3k +2)

OpenStudy (anonymous):

I have that much so far. Basically I am trying to get that to reduce to k(k+1)(k+2) plus whatever is left which 3 should divide into. k(k+1)(k+2) is already assumed divide by 3 that

OpenStudy (anonymous):

(k+1)(k+2)(k+3)=1/3 (3k+1)(k+2) (k+1) proved

OpenStudy (anonymous):

I'm not really following that.... How did you get that?

OpenStudy (anonymous):

you want to show in third step that it is divisible by 3 that I have shown you

OpenStudy (anonymous):

if you now now multiply (3k+1)(k+2)(k+1), whatever you will get it must be divisible by 3

OpenStudy (anonymous):

because 1/3 is now thier factor

OpenStudy (anonymous):

Ok, I see. I wasn't sure where the 1/3 came from. Thanks